if the perimeter of a rectangle is 24 and the height and the length of the base are in a ratio of 3:1, what are the dimensions of the rectangle
Answer by jim_thompson5910(35256)

A ratio of 3:1 means that width is x and the length is 3x

P = 2L + 2W

24 = 2(3x) + 2x

24 = 6x + 2x

24 = 8x

24/8 = 8x/8

3 = x

x = 3


So the width is 3 and the length is 3*3 = 9